# SAGE: Scalable AI Governance & Evaluation

### v1

@@ 4.1. Implementation Decisions & Trade-offs
Our primary design objective is governance utility, prioritizing rapid iteration, interpretability, and actionable diagnostics over maximizing model sophistication. To this end, we explicitly restricted the Judge’s scope to semantic relevance, decoupling it from mathematical signals like network distance or job posting recency which can be calculated deterministically.
@@
Trade-off: We intentionally accepted a smaller, expert-curated Precedent set to prioritize label accuracy over quantity, ensuring policy agility. A compact, expert-maintained precedent enables rapid realignment: after a policy revision, the Precedent can be updated within hours, keeping the model synchronized with evolving product judgment.
@@ 4.2. The Bidirectional Calibration Loop
Judge $\rightarrow$ Precedent (Adversarial Audit): We recognize that Precedent is susceptible to human error. When the Judge disagrees with a human label, we audit the reasoning trace to understand the model’s decision and confirm whether the policy was applied accurately. We found that the Judge frequently correctly identified evidence buried in verbose documents (e.g., detailed job postings or lengthy member profiles) that human reviewers overlooked. Upon expert validation, we update the Precedent, leveraging the model to surface potential errors for human correction.
@@
Table 1 shows the efficacy of these feedback vectors as Judge-Precedent agreement improves with each intervention. For Semantic People Search, through several iterations of bidirectional calibration, we achieved Teacher–Human agreement of 0.77 (linear weighted Cohen’s kappa), surpassing the 0.7 threshold for substantial agreement (Landis and Koch, 1977) and approaching the practical agreement ceiling of expert product owners (0.83; Figure 3) in a complex and often subjective domain, establishing a stable foundation for distillation.
@@ 5.2. Distillation at Scale
We train the Student Judge via supervised instruction tuning (Wei et al., 2021) over hydrated prompts paired with Teacher Judge outputs. While early iterations employed parameter-efficient LoRA (Hu et al., 2021), full-parameter fine-tuning delivered significantly higher agreement stability for graded judgments. The final production Student therefore adopts full fine-tuning with ZeRO-3–style optimizer (Rajbhandari et al., 2020) and state partitioning, sharding model parameters, optimizer states, and gradients across workers for efficient training.
@@ 5.3. Experiments: Optimizing the Student Model
Figure 3 further contextualizes alignment ceilings and distillation fidelity. Human–Human agreement (0.78 for Job Search; 0.83 for People Search) represents the practical ceiling of this task: even expert annotators do not fully agree, so no judge should be expected to exceed this limit in a meaningful way. The Teacher–Human alignment (0.76 for Job Search; 0.77 for People Search) forms the effective upper bound for distillation: since the Student is trained to imitate the Teacher, the Student–Human alignment is naturally constrained by how well the Teacher itself matches human precedent. In this setting, the Student achieves 0.72 for Job Search; 0.73 for People Search, approaching the Teacher’s ceiling while operating at production scale. Additionally, as an illustrative example in People Search, Student-Teacher agreement reaches 0.81, indicating faithful distillation without significant loss of judgment behavior.
Figure 2. Student Judge linear kappa across major training iterations for Job Search and People Search.
@@
Figure 3. Linear weighted Cohen’s kappa scores across different judge comparisons.
@@ 6.4. Online Experimentation
Figure 5. PMR@10 vs. Dismiss to Apply Ratio (D2A) in the presence of a production incident degrading recommendation quality.
@@
    Sampling Strategy: We choose to stratify-sample along different dimensions for offline candidate selection vs. A/B experiment decisions. For offline candidate selection, the focus is typically on upsampling and improving system performance on long-tail queries that are perhaps less common (e.g., foreign language). Meanwhile, during online experimentation the primary focus is high-resolution measurement across all of the major product channels/surfaces where we serve recommendations (e.g., email, jobs home, feed). This not only reflects how we structure development of our system, but also ensures that excessively high-volume channels don’t bias our understanding of quality throughout the ecosystem.
Once a winning variant is chosen and ramped as the Majority Member Experience (MME), we continue to surface policy-based relevance metrics derived from the daily pipeline in operational dashboards to monitor system health on an ongoing basis. This enables us to catch system regressions that don’t immediately manifest as declines in engagement. For example, Figure 5 shows a production incident where recommendation quality degraded, causing a 40% week-over-week increase in PMR@10, yet our closest engagement-based proxy for relevance, Dismiss to Apply Ratio (D2A), did not meaningfully deviate from neutral.

### v2

@@ 4.1. Implementation Decisions & Trade-offs
Our primary design objective is governance utility, prioritizing rapid iteration, interpretability, and actionable diagnostics over maximizing model expressiveness. To this end, we explicitly restricted the Judge’s scope to semantic relevance, decoupling it from mathematical signals like network distance or job posting recency which can be calculated deterministically.
@@
Trade-off: We intentionally accepted a smaller, expert-curated Precedent set to prioritize label accuracy over quantity, thereby enabling policy agility. A compact, expert-maintained precedent enables rapid realignment: after a policy revision, the Precedent can be updated within hours, keeping the model synchronized with evolving product judgment.
@@ 4.2. The Bidirectional Calibration Loop
Judge $\rightarrow$ Precedent (Adversarial Audit): We recognize that Precedent is susceptible to human error. When the Judge disagrees with a human label, we audit the reasoning trace to understand the model’s decision and confirm whether the policy was applied accurately. We found that the Judge frequently correctly identified evidence buried in verbose documents (e.g., detailed job postings or lengthy member profiles) that human reviewers overlooked. Upon expert validation, we update the Precedent, leveraging the Judge to surface potential errors for human correction.
@@
Table 1 shows the efficacy of these feedback vectors as Judge-Precedent agreement improves with each intervention. For Semantic People Search, through several iterations of bidirectional calibration, we achieved Teacher Judge–Human agreement of 0.77 (linear weighted Cohen’s kappa), surpassing the 0.7 threshold for substantial agreement (Landis and Koch, 1977) and approaching the practical agreement ceiling of expert product owners (0.83; Figure 3) in a complex and often subjective domain, establishing a stable foundation for distillation.
@@ 5.2. Distillation at Scale
We train the Student Judge via supervised instruction tuning (Wei et al., 2021) over hydrated prompts paired with Teacher Judge outputs. While early iterations employed parameter-efficient LoRA (Hu et al., 2021), we found that full-parameter fine-tuning delivered significantly higher agreement stability for graded judgments. The final production Student therefore adopts full-parameter fine-tuning with ZeRO-3–style optimizer (Rajbhandari et al., 2020). By sharding model parameters, gradients, and optimizer states across workers, this approach ensures memory-efficient training at scale.
@@ 5.3. Experiments: Optimizing the Student Model
Figure 3 further contextualizes alignment ceilings and distillation fidelity. Human–Human agreement (0.78 for Job Search; 0.83 for People Search) represents the practical ceiling of this task: even expert annotators do not fully agree, so no judge should be expected to exceed this limit in a meaningful way. The Teacher–Human alignment (0.76 for Job Search; 0.77 for People Search) forms the effective upper bound for distillation: since the Student is trained to imitate the Teacher, the Student–Human alignment is naturally constrained by how well the Teacher itself matches human precedent. In this setting, the Student achieves 0.72 for Job Search; 0.73 for People Search, approaching the Teacher’s ceiling while operating at production scale. Additionally, as an illustrative example in People Search, Student-Teacher agreement reaches 0.81, indicating faithful transfer of reasoning logic.
Figure 2. Linear Cohen’s kappa of Student Judge across major training iterations for Job Search and People Search.
@@
Figure 3. Linear Cohen’s kappa scores across different judge comparisons.
@@ 6.4. Online Experimentation
Figure 5. PMR@10 vs. Dismiss to Apply Ratio (D2A) in the presence of a production incident degrading search quality.
@@
    Sampling Strategy: We choose to stratify-sample along different dimensions for offline candidate selection vs. A/B experiment decisions. For offline candidate selection, the focus is typically on upsampling and improving system performance on long-tail queries that are perhaps less common (e.g., foreign language). Meanwhile, during online experimentation the primary focus is high-resolution measurement across all of the major product channels/surfaces (e.g., email, jobs home, feed). This not only reflects how we structure development of our system, but also ensures that excessively high-volume channels don’t bias our understanding of quality throughout the ecosystem.
Once a winning variant is chosen and ramped as the Majority Member Experience (MME), we continue to surface policy-based relevance metrics derived from the daily pipeline in operational dashboards to monitor system health on an ongoing basis. This enables us to catch system regressions that don’t immediately manifest as declines in engagement. For example, Figure 5 shows a production incident where search quality degraded, causing a 40% week-over-week increase in PMR@10, yet our closest engagement-based proxy for relevance, Dismiss to Apply Ratio (D2A), did not meaningfully deviate from neutral.